# Treewidth of generalized Hamming graph, bipartite Kneser graph and generalized Petersen graph

### v1

@@ 3.2 Treewidth of $H(t,q,n)$
In this subsection we analyze the asymptotic behavior of $tw(H(t,q,n))$ when $n$ goes to infinity and give proof of Theorem 1.2.
@@
$tw(H(t,q,n))\geq c_{1}tq^{n}/\sqrt{n}$ for some constant $c_{1}$.
@@
for some constant $c_{1}>0$. Then, $b_{v}(m,H(t,q,n))\geq c_{1}tq^{n}/\sqrt{n}$. $\square$
@@
$tw(H(t,q,n))\leq pw(H(t,q,n))\leq c_{2}tq^{n}/\sqrt{n}$ for some constant $c_{2}$.
@@
$f(i)=\left\{\begin{array}[]{cc}0&\mbox{if $1\leq i\leq q/2$,}\\ 1&\mbox{if $q/2<i\leq q$.}\\ \end{array}\right.$
@@
Combining Lemmas 3.15 and 3.16, we can derive Theorem 1.2.
@@ 4 Treewidth of bipartite Kneser graph and Johnson graph
For positive integers $n$ and $k$ satisfying $n\geq 2k+1$, the bipartite Kneser graph $BK(n,k)$ has all subsets of $[n]$ with $k$ or $n-k$ elements as vertices and an edge between any two vertices when one is a subset of the other. It is also called middle cube graph. In the following, we will use $k$-subsets and $(n-k)$-subsets of $[n]$ to represent the vertices of bipartite Kneser graph. We call vertices that are $k$-subsets of $[n]$ the left part denoted by $V_{L}$, and the rest is called the right part denoted by $V_{R}$. $V_{L}$ and $V_{R}$ are two parts of bipartite Kneser graph $BK(n,k)$.
For positive integers $n$ and $k$ satisfying $n>k$, the Johnson graph $J(n,k)$ has all subsets of $[n]$ with $k$ elements and an edge between any two vertices when their intersection has exactly $(k-1)$ elements. Since we have a bijection between all $k$-subsets of $[n]$ and all binary $n$-vectors with exactly $k$ ones, we can also treat a vertex of $J(n,k)$ as a $n$-vector with exactly $k$ ones. From this point of view, two vertices are adjacent iff their the Hamming distance of their corresponding $n$-vectors is no more than $2$. Therefore, $J(n,k)$ is the $k$-th slice of $H(2,2,n)$, that is, $J(n,k)$ is a subgraph of $H(2,2,n)$ induced by vertices from $H(2,2,n)$ corresponding to $n$-vector with exactly $k$ ones.
@@ 4.1 Treewidth of $BK(n,k)$ when $n$ is large enough
In this section, we focus on the treewidth of $BK(n,k)$ when $n$ is large enough and give proof of Theorem 1.3. Before proof, we need the following proposition.
@@
Proof.  Denote $BK(n,k)$ by $G$ and $V(G)=V_{L}\cup V_{R}$, where $V_{L}=\binom{[n]}{k}$ and $V_{R}=\binom{[n]}{n-k}$. Suppose $tw(G)<\binom{n}{k}-1$. From Corollary 2.5, there exists a separator $X$ of $G$ with $|X|<\binom{n}{k}$ such that there exists non-empty vertex set $A$ and $B$ with $A\cup B=V(G)-X$, $A\cap B=\emptyset$, $|V(G)-X|/3\leq|A|,|B|\leq 2|V(G)-X|/3$ and there is no edge between $A$ and $B$. Let $A_{L}=A\cap V_{L}$, $A_{R}=A\cap V_{R}$, $B_{L}=B\cap V_{L}$ and $B_{R}=B\cap V_{R}$.
@@
$\mathcal{A}$ =$=\left\{S\in\binom{[n]}{k}\mid S\in A_{L}\right\},$
$\mathcal{B}$ =$=\left\{S\in\binom{[n]}{n-k}\mid S\in B_{R}\right\}.$
Let $\mathcal{C}=\{[n]-S\mid S\in\mathcal{B}\}$. Since there is no edge between $A_{L}$ and $B_{R}$, for any $S_{1}\in\mathcal{A}$ and $S_{2}\in\mathcal{B}$, $S_{1}\nsubseteq S_{2}$ which implies $S_{1}\cap([n]-S_{2})\neq\emptyset$. Hence, $\mathcal{A}$ and $\mathcal{C}$ is cross-intersecting. From [11], we have $|\mathcal{A}|+|\mathcal{C}|\leq\binom{n}{k}-\binom{n-k}{k}+1$. Notice that $|\mathcal{A}|=|A_{L}|$ and $|\mathcal{C}|=|\mathcal{B}|=|B_{R}|$. Then we have $|A_{L}|+|B_{R}|\leq\binom{n}{k}-\binom{n-k}{k}+1$.
@@
If there is only one empty set in $\{A_{R},B_{L}\}$, say $A_{R}\neq\emptyset$ and $B_{L}=\emptyset$, then we have $|B|=|B_{R}|\leq\binom{n}{k}-\binom{n-k}{k}$ by $|A_{L}|+|B_{R}|\leq\binom{n}{k}-\binom{n-k}{k}+1$ and $A_{L}\not=\emptyset$. Since $|V(G)-X|/3\leq|A|,|B|\leq 2|V(G)-X|/3$, we have $|A|+|B|\leq 3|B|\leq 3\left(\binom{n}{k}-\binom{n-k}{k}\right)$. Since $3\binom{n-k}{k}\geq 2\binom{n}{k}$, we have $|X|=|V(G)|-|A|-|B|\geq 3\binom{n-k}{k}-\binom{n}{k}\geq\binom{n}{k}$, a contradiction. $\square$
Theorem 1.3 can be easily derived from Lemmas 4.3 and 4.4.
@@ 4.2 Treewidth of $BK(2k+1,k)$ and $J(2k+1,k)$
When $n$ is large enough, the treewidth of bipartite Kneser graph can be exactly calculated by Theorem 1.3. Now we focus on the treewidth of bipartite Kneser graph when $n$ is small and give proof of Theorem 1.4.
In order to prove our reslt, we need more definitions. A graph $G$ is chordal if and only if, in any cycle of length larger than 3 in $G$, there exists a chord connecting two nonadjacent vertices of the cycle. Given a graph $G$, define $\omega(G)$, the clique number of $G$ to be the number of vertices of the largest clique in $G$.
@@
Proof.Let $n=2k+1$. By Proposition 4.5, there exists a chordal graph $H$ such that $J(n,k)\subseteq H$ and $\omega(H)-1=tw(J(n,k))$. Add edges to the left part of $BK(n,k)$ such that the left part is isomorphic to $H$ and denote the result graph as $BK^{\prime}(n,k)$. That is, $BK^{\prime}(n,k)[V_{L}]\cong H$.
@@
Lemma 4.10 can be derived from Propositions 4.8 and 4.9. Theorem 1.4 can be derived by Lemmas 4.64.7 and 4.10.
@@ 5 Treewidth of generalized Petersen graph
$V(G_{n,k})$ =$=\{v_{1},\ldots,v_{n},u_{1},\ldots,u_{n}\},$
$E(G_{n,k})$ =$={\{v_{i}u_{i}\}}\cup{\{v_{i}v_{i+1}\}}\cup{\{u_{i}u_{i+k}\}},i=1,2,\ldots,n,$
@@
With the help of Proposition 5.1, we now can give proof of Theorem 1.5.
@@
  • If $i+t+tk<j<n+i-t$, then $j+t<i+n<j+n-t-tk\leq j+t+tk$. The last inequality comes from $2kt+2t\geq n$ by $t=\lceil\frac{n}{2k+2}\rceil$ and then we have $t+tk\geq n-t-tk$ which implies $j+t<i+n\leq j+t+tk$. The next proof is the same as that in the second situation.
@@
Then $B$ is a bramble. Let $S$ be a hitting set of $B$. We construct a hypergraph $H$ with vertex set $V(H)=V(G_{n,k})$ and hyperedge set ${\{V_{i}\}}_{1\leq i\leq n}$. Then $S$ is a transversal of $H$ and thus, $\min|S|=\tau(H)$ where $\tau(H)$ is the transversal number of $H$. Since $\tau(H)\geq\max\limits_{H^{\prime}\subseteq H}\frac{m(H^{\prime})}{\Delta(H^{\prime})}$ [3], the order of $B$ is
@@
$2k+2-\frac{n}{t+1}\leq\frac{n}{t-1}-\frac{n}{t+1}=\frac{2n}{t^{2}-1}\leq 1$
Therefore, the order of $B$ is no less than $22k+$ 2. From Proposition 5.1, we can derive that $tw(G_{n,k})\geq 2k+1$.
@@
Since $|X_{i}|=|Z_{i}|=k+1,|Y_{i}|=|W_{i}|=k+2,|B_{1}|=k,|B_{2}|=k+1$ and $|A|=k+1$, the width of $\mathcal{P}_{2}$ is $32k+$ 3 and, hence, $tw(G_{n,k})\leq pw(G_{n,k})\leq 2k+2$. $\square$
@@ Acknowledgement
M. Cao is supported by the National Natural Science Foundation of China (Grant 12301431), M. Lu is supported by the National Natural Science Foundation of China (Grant 12171272 & 12161141003).
@@
First, it is easy to verify that $D_{t,n,k,s}=A^{(1)}_{t,n,k,s}+A^{(2)}_{t,n,k,s}+C_{t,n,k,s}$ holds when $k-s=0$ and, $D_{t,n,k,s}=B_{t,n,k,s}+C_{t,n,k,s}$ when $k-s=n-t$ and $A^{(1)}_{t,n,k,s}+A^{(2)}_{t,n,k,s}=B_{t,n,k,s}$ hold when $k-s=\lfloor(n-t)/2\rfloor$. Hence, in those situations, we can calculate $r(M^{(t,n)}_{k,k+t-2s})$ in both ways.
@@
When $t=1$, then $s=0$. In this case $r^{(2)}_{1,n,k,1}=-\infty$ and $r^{(3)}_{1,n,k,1}=-\infty$ and, hence, $r(M_{k,k+1}^{(1,n)})=\max\{r^{(1)}_{1,n,k,1},r^{(4)}_{1,n,k,1}\}$. We can prove Eq 13 holds for $t=1$ by induction on $n$. The proof is omitted.
Then, we give a proof of Lemma 3.10 via Eq 12 by induction. Suppose Eq 13 holds for $t+n+k+s\leq N-1$. Now consider when $t+n+k+s=N$. In the following steps, we only need to consider the case $t\geq 2$ and $t<n-1$.
Step 2: verify Eq 13 when $k=0$.
If $k=0$, then $k+t-2s<n$ and $r^{(1)}_{t,n,k,t-2s}=-\infty$ and $r^{(2)}_{t,n,k,t-2s}=-\infty$.
@@
Case 2.2: If $t>2s$, then $(t-1,n-1,k,s)$ are valid and, hence, $r^{(3)}_{t,n,k,t-2s}=r(M^{(t-1,n-1)}_{k,k+(t-1)-2s})$ can be calculated via Eq 13 by induction.
@@
Subcase 2.2.2: If $t<n-2$, then $(t,n-1,k,s)$ are valid and, hence, $r^{(4)}_{t,n,k,t-2s}=r(M^{(t,n-1)}_{k,k+t-2s})$ can be calculated via Eq 13 by induction.
@@
Step 3: verify Eq 13 when $k+t-2s=n$.
In this case, $r^{(2)}_{t,n,k,t-2s}=-\infty$ and $r^{(4)}_{t,n,k,t-2s}=-\infty$.
@@
Case 3.2: If $t<2s$, then $(t-1,n-1,k,s)$ is valid and, hence, $r_{t,n,k,p}^{(3)}=r(M_{k,k+p-1}^{(t-1,n-1)})$ can be calculated via Eq 13 by induction. Since $k-s\geq n-t$, we have
@@
$=\binom{n-1}{k-1}+\binom{n-1}{k-1+t-2s}-1+\binom{n-1}{k}$
$=\binom{n}{k}+\sum_{m=k+t-2s+1}^{n}\binom{m-1}{k+t-2s-1}=\binom{n}{k}+D_{t,n,k,s},$
where the second equality comes from Step 1 and the following equalities come from $k+t-2s=n$.
Subcase 3.2.1: If $t<n-2$, then $(t,n-1,k-1,s)$ is valid and, hence, $r^{(1)}_{t,n,k,t-2s}=r(M^{(t,n-1)}_{k-1,k-1+t-2s})+\binom{n-1}{k}$ can be calculated via Eq 13 by induction. That is
@@
where the second equality comes from Eq 13 by induction, and the third equality comes from $k+t-2s=n$.
From above, combining Eq 12, we can verify that Eq 13 holds when $k+t-2s=n$.
Step 4: verify Eq 13 when $k>0$ and $k+t-2s<n$.
@@
Note that $(t-1,n-1,k-1,s-1)$ is valid and, hence, we can calculate $r^{(2)}_{t,n,k,t-2s}$ via Eq 13 by induction.
Case 4.1.1: If $k-s<0$ or $k-s>n-t$, then $(k-1)-(s-1)<0$ or $(k-1)-(s-1)>(n-1)-(t-1)$.
@@
$\quad+\sum\limits_{m=t-3s+1+2k}^{n-s}\left(\binom{m-1}{k+t-2s-1}-\binom{m-1}{k-s-1}\right)+\sum\limits_{m=n-s+1}^{n}\binom{m-1}{k+t-2s-1}$
@@
Case 4.2.1: If $(k-1)-s<0$, then $(k-1)-(s-1)=k-s\leq 0$ and
@@
Case 4.2.2: If $0\leq(k-1)-s<\lfloor((n-1)-t)/2\rfloor$, then $0\leq 1\leq k-s\leq\lfloor(n-t-1)/2\rfloor\leq\lfloor(n-t)/2\rfloor$ and
$A^{(1)}_{t,n-1,k-1,s}-A^{(1)}_{t-1,n-1,k-1,s-1}=-\left(\binom{t-2k+s+2}{t+k-2s-2}-\binom{t-2k+s+2}{k-s-2}\right),$
@@
$=$ $(B_{t,n-1,k-1,s}-B_{t-1,n-1,k-1,s-1})+(C_{t,n-1,k-1,s}-C_{t-1,n-1,k-1,s-1})$
$=$ $C_{t,n-1,k-1,s}-C_{t-1,n-1,k-1,s-1}$
$=$ $\binom{n-2}{k+t-2s-2}-\binom{n-1}{k-1+t-2s}-\sum_{m=n-s}^{n-1}\binom{m-1}{k+t-2s-1}$
$\leq$ $0$
Case 4.2.4: If $(k-1)-s>(n-1)-t$, then $k-s>n-t$.

### v3

@@ 3.2 Treewidth of $H(t,q,n)$
In this subsection we analyze the asymptotic behavior of $tw(H(t,q,n))$ when $n$ goes to infinity and give the proof of Theorem 1.2.
@@
$tw(H(t,q,n))\geq c_{1}tq^{n}/\sqrt{n}$ for some constant $c_{1}$ not depending on $t$ or $q$ when $n$ is sufficiently large.
@@
for some constant $c_{1}>0$ not depending on $t$ and $q$. Then, $b_{v}(m,H(t,q,n))\geq c_{1}tq^{n}/\sqrt{n}$. $\square$
@@
$tw(H(t,q,n))\leq pw(H(t,q,n))\leq c_{2}tq^{n}/\sqrt{n}$ for some constant $c_{2}$ not depending on $t$ or $q$ when $n$ is sufficiently large.
@@
[…]
@@
Combining Lemmas 3.16 and 3.17, we can derive Theorem 1.2.
@@ 4 Treewidth of bipartite Kneser graph and Johnson graph
For positive integers $n$ and $k$ satisfying $n\geq 2k+1$, the bipartite Kneser graph $BK(n,k)$ has all subsets of $[n]$ with $k$ or nkn-k elements as vertices and an edge between any two vertices when one is a subset of the other. It is also called middle cube graph. In the following, we will use $k$-subsets and (nk)(n-k)-subsets of $[n]$ to represent the vertices of bipartite Kneser graph. We call vertices that are $k$-subsets of $[n]$ the left part denoted by $V_{L}$, and the rest is called the right part denoted by $V_{R}$. $V_{L}$ and $V_{R}$ are two parts of bipartite Kneser graph $BK(n,k)$.
For positive integers $n$ and $k$ satisfying $n>k$, the Johnson graph $J(n,k)$ has all subsets of $[n]$ with $k$ elements and an edge between any two vertices when their intersection has exactly (k1)(k-1) elements. Since we have a bijection between all $k$-subsets of $[n]$ and all binary $n$-vectors with exactly $k$ ones, we can also treat a vertex of $J(n,k)$ as an $n$-vector with exactly $k$ ones. From this point of view, two vertices are adjacent iff their the Hamming distance of their corresponding $n$-vectors is no more than $2$. Therefore, $J(n,k)$ is the $k$-th slice of $H(2,2,n)$, that is, $J(n,k)$ is a subgraph of $H(2,2,n)$ induced by vertices from $H(2,2,n)$ corresponding to $n$-vector with exactly $k$ ones.
@@ 4.1 Treewidth of $BK(n,k)$ when $n$ is large enough
In this section, we focus on the treewidth of $BK(n,k)$ when $n$ is large enough and give the proof of Theorem 1.3. Before proof, we need the following proposition.
@@
Proof.    Denote $BK(n,k)$ by $G$ and $V(G)=V_{L}\cup V_{R}$, where $V_{L}=\binom{[n]}{k}$ and $V_{R}=\binom{[n]}{n-k}$. Suppose $tw(G)<\binom{n}{k}-1$. From Corollary 2.5, there exists a separator $X$ of $G$ with $|X|<\binom{n}{k}$ such that there exists non-empty vertex set $A$ and $B$ with $A\cup B=V(G)-X$, $A\cap B=\emptyset$, $|V(G)-X|/3\leq|A|,|B|\leq 2|V(G)-X|/3$ and there is no edge between $A$ and $B$. Let $A_{L}=A\cap V_{L}$, $A_{R}=A\cap V_{R}$, $B_{L}=B\cap V_{L}$ and $B_{R}=B\cap V_{R}$ (see Figure 3).
Figure 3: Sketch graph of $BK(n,k)$. The condition that there is no edge between $A$ and $B$ equals to there is no edge between $A_{L}$ and $B_{R}$, and between $A_{R}$ and $B_{R}$.
@@
$\mathcal{A}$ $=\left\{S\in\binom{[n]}{k}\mid S\in A_{L}\right\},$
$\mathcal{B}$ $=\left\{S\in\binom{[n]}{n-k}\mid S\in B_{R}\right\}.$
Let $\mathcal{C}=\{[n]-S\mid S\in\mathcal{B}\}\subseteq\binom{[n]}{k}$. Since there is no edge between $A_{L}$ and $B_{R}$, for any $S_{1}\in\mathcal{A}$ and $S_{2}\in\mathcal{B}$, $S_{1}\nsubseteq S_{2}$ which implies $S_{1}\cap([n]-S_{2})\neq\emptyset$.
Definition 4.5 (Cross-intersecting families)
Let $\mathcal{A}$ and $\mathcal{B}$ be two families of subsets of a finite set $X$. We say that $\mathcal{A}$ and $\mathcal{B}$ are cross-intersecting if for any $A\in\mathcal{A}$ and $B\in\mathcal{B}$, $A\cap B\neq\emptyset$.
Hence, $\mathcal{A}$ and $\mathcal{C}$ is cross-intersecting. By the properties of cross-intersecting families from [13], we have $|\mathcal{A}|+|\mathcal{C}|\leq\binom{n}{k}-\binom{n-k}{k}+1$. Notice that $|\mathcal{A}|=|A_{L}|$ and $|\mathcal{C}|=|\mathcal{B}|=|B_{R}|$. Then we have $|A_{L}|+|B_{R}|\leq\binom{n}{k}-\binom{n-k}{k}+1$.
@@
If there is only one empty set in $\{A_{R},B_{L}\}$, say $A_{R}\neq\emptyset$ and $B_{L}=\emptyset$, then we have $|B|=|B_{R}|\leq\binom{n}{k}-\binom{n-k}{k}$ by $|A_{L}|+|B_{R}|\leq\binom{n}{k}-\binom{n-k}{k}+1$ and $A_{L}\not$$\emptyset$. Since $|V(G)-X|/3\leq|A|,|B|\leq 2|V(G)-X|/3$, we have $|A|+|B|\leq 3|B|\leq 3\left(\binom{n}{k}-\binom{n-k}{k}\right)$. Since $3\binom{n-k}{k}\geq 2\binom{n}{k}$, we have $|X|=|V(G)|-|A|-|B|\geq 3\binom{n-k}{k}-\binom{n}{k}\geq\binom{n}{k}$, a contradiction. $\square$
Theorem 1.3 can be easily derived from Lemmas 4.3 and $4.$
@@ 4.2 Treewidth of $BK(2k+1,k)$ and $J(2k+1,k)$
When $n$ is large enough, the treewidth of the bipartite Kneser graph can be exactly calculated by Theorem 1.3. Now we focus on the treewidth of the bipartite Kneser graph when $n$ is small and give the proof of Theorem 1.4.
In order to prove our reslt, we need more definitions. A graph $G$ is chordal if and only if, in any cycle of length larger than 3 in $G$, there exists a chord connecting two nonadjacent vertices of the cycle. Given a graph $G$, define $\omega(G)$, the clique number of $G$ to be the number of vertices of the largest clique in $G$. The treewidth of a graph $G$ has a close relationship with its chordal supergraph as Proposition 4.6 shows.
@@
Using Proposition 4.6, we can prove the relationship between the treewidth of the bipartite Kneser graph and Johnson graph as Lemma 4.7. We first explain the proof ideas. Note that $BK(n,k)$ is a bipartite graph with two parts $\binom{[n]}{k}$ and $\binom{[n]}{n-k}$. And the vertex set of $J(n,k)$ is $\binom{[n]}{k}$ which is the same as one part of $BK(n,k)$. Thus, we can embed $J(n,k)$ into $BK(n,k)$ by letting the $\binom{[n]}{k}$ part of $BK(n,k)$ is isomorphic to $J(n,k)$. We then will show the result graph is chordal. After calculating its clique number and then using Proposition 4.6, we can derive our result.
@@
Proof.    Let $n=2k+1$. By Proposition 4.6, there exists a chordal graph $H$ such that $J(n,k)$ is a subgraph of $H$ and $\omega(H)-1=tw(J(n,k))$. Add edges to the left part of $BK(n,k)$ such that the left part is isomorphic to $H$ and denote the result graph as $BK^{\prime}(n,k)$. That is, $BK^{\prime}(n,k)[V_{L}]\cong H$.
@@
Lemma 4.11 can be derived from Propositions 4.9 and 4.10. Theorem 1.4 can be derived by Lemmas 4.74.8 and 4.11.
@@ 5 Treewidth of generalized Petersen graph
$V(G_{n,k})$ $=\{v_{1},\ldots,v_{n},u_{1},\ldots,u_{n}\},$
$E(G_{n,k})$ $={\{v_{i}u_{i}\}}\cup{\{v_{i}v_{i+1}\}}\cup{\{u_{i}u_{i+k}\}},i=1,2,\ldots,n,$
@@
With the help of Proposition 5.1, we now can give the proof of Theorem 1.5.
@@
  • If i+t+tk<j<n+iti+t+tk<j<n+i-t, then $j+t<i+n<j+n-t-tk\leq j+t+tk$. The last inequality comes from $2kt+2t\geq n$ by $t=\lceil\frac{n}{2k+2}\rceil$ and then we have $t+tk\geq n-t-tk$ which implies $j+t<i+n\leq j+t+tk$. The next proof is the same as that in the second situation.
@@
Then $B$ is a bramble. Let $S$ be a hitting set of $B$. We construct a hypergraph $H$ with vertex set $V(H)=V(G_{n,k})$ and hyperedge set ${\{V_{i}\}}_{1\leq i\leq n}$.
Definition 5.2 (Transversal)
Let $H$ be a hypergraph on a set $X$ with edges $E_{1},\ldots,E_{m}$. A set $T\subseteq X$ is called a transversal of $H$ if $T$ intersects every edge of $H$, that is,
$T\cap E_{i}\neq\emptyset,\forall i=1,2,\ldots,m.$
Then $S$ is a transversal of $H$ and thus, $\min|S|=\tau(H)$ where $\tau(H)$ is the transversal number of $H$. Since $\tau(H)\geq\max\limits_{H^{\prime}\subseteq H}\frac{m(H^{\prime})}{\Delta(H^{\prime})}$ [4], where $m(H^{\prime})$ is the number of edges in $H^{\prime}$ and $\Delta(H^{\prime})$ is the maximum degree of $H^{\prime}$, we have that the order of $B$ is
@@
$2k+2-\frac{n}{t+1}\leq\frac{n}{t-1}-\frac{n}{t+1}=\frac{2n}{t^{2}-1}<1.$
Therefore, the order of $B$ is no less than $2k+2$. From Proposition 5.1, we can derive that $tw(G_{n,k})\geq 2k+1$.
@@
Here we explain why we construct the path-decomposition in this way. For a cycle $v_{1}v_{2}\ldots v_{k}$, the way to build its path-decomposition is to first delete any vertex, say $v_{1}$, then the left part is a path $v_{2}\ldots v_{k}$. Then build a path-decomposition like $\{v_{2}v_{3}\}-\{v_{3}v_{4}\}-\ldots-\{v_{k-1}v_{k}\}$ and then add $v_{1}$ to all bags. Here the generalized Petersen graph behaves like a “double-cycle”. We first delete $A$, and the rest part behaves like a “path”. Then the sequence of $X_{i},Y_{i},Z_{i},W_{i},X_{i+1}$ is like what we do in the path-decomposition of a cycle. Once we move one vertex so that the width do not increase too much. $B_{1}$ and $B_{2}$ are designed to cover the left vertices. Finally we add $A$ to all bags just like for the cycle we add $v_{1}$ back.
@@
Since $|X_{i}|=|Z_{i}|=k+1,|Y_{i}|=|W_{i}|=k+2,|B_{1}|=k,|B_{2}|=k+1$ and $|A|=k+1$, the width of $\mathcal{P}_{2}$ is $2k+3$ and, hence, $tw(G_{n,k})\leq pw(G_{n,k})\leq 2k+2$. $\square$
@@ Acknowledgement
The authors would like to thank the referee for their careful reading and valuable comments which help to improve the presentation of this paper. M. Cao is supported by the National Natural Science Foundation of China (Grant 12301431), M. Lu is supported by the National Natural Science Foundation of China (Grant 12171272).
@@
First, it is easy to verify that $D_{t,n,k,s}=A^{(1)}_{t,n,k,s}+A^{(2)}_{t,n,k,s}+C_{t,n,k,s}$ holds when ks=0k-s=0 and, $D_{t,n,k,s}=B_{t,n,k,s}+C_{t,n,k,s}$ when ks=ntk-s=n-t and $A^{(1)}_{t,n,k,s}+A^{(2)}_{t,n,k,s}=B_{t,n,k,s}$ hold when $k-s=\lfloor(n-t)/2\rfloor$. Hence, in those situations, we can calculate $r(M^{(t,n)}_{k,k+t-2s})$ in both ways.
In the following, we prove the Lemma by induction on $t+n+k+s$. From Table 1, when we calculate $\{r^{(i)}_{t,n,k,t-2s}\}_{1\leq i\leq 4}$, the term always obtains a smaller $t^{\prime}+n^{\prime}+k^{\prime}+s^{\prime}$, which means we can use Eq 13 to calculate those terms by induction.
@@
When $t=1$, then $s=0$. In this case $r^{(2)}_{1,n,k,1}$$\infty$ and $r^{(3)}_{1,n,k,1}$$\infty$ and, hence, $r(M_{k,k+1}^{(1,n)})=\max\{r^{(1)}_{1,n,k,1},r^{(4)}_{1,n,k,1}\}$. We can prove Eq 13 holds for $t=1$ by induction on $n$. The proof is omitted.
Then, we give a proof of Lemma 3.11 via Eq 12 by induction. Suppose Eq 13 holds for $t+n+k+s\leq N-1$. Now consider when $t+n+k+s=N$. In the following steps, we only need to consider the case $t\geq 2$ and t<n1t<n-1.
Step 2: verify Eq 13 when $k=0$ and 1<t<n11<t<n-1.
If $k=0$, then k+t2s<nk+t-2s<n and $r^{(1)}_{t,n,k,t-2s}$$\infty$ and $r^{(2)}_{t,n,k,t-2s}$$\infty$.
@@
Case 2.2: If $t>2s$, then (t1,n1,k,s)(t-1,n-1,k,s) are valid and, hence, $r^{(3)}_{t,n,k,t-2s}=r(M^{(t-1,n-1)}_{k,k+(t-1)-2s})$ can be calculated via Eq 13 by induction.
@@
Subcase 2.2.2: If t<n2t<n-2, then (t,n1,k,s)(t,n-1,k,s) are valid and, hence, $r^{(4)}_{t,n,k,t-2s}=r(M^{(t,n-1)}_{k,k+t-2s})$ can be calculated via Eq 13 by induction.
@@
Step 3: verify Eq 13 when k+t2s=nk+t-2s=n and 1<t<n11<t<n-1.
In this case, $r^{(2)}_{t,n,k,t-2s}$$\infty$ and $r^{(4)}_{t,n,k,t-2s}$$\infty$.
@@
Case 3.2: If $t<2s$, then (t1,n1,k,s)(t-1,n-1,k,s) is valid and, hence, $r_{t,n,k,p}^{(3)}=r(M_{k,k+p-1}^{(t-1,n-1)})$ can be calculated via Eq 13 by induction. Since $k-s\geq n-t$, we have
@@
$=\binom{n-1}{k}+\binom{n-1}{k-1}+\binom{n-1}{k-1+t-2s}-1$
$=\binom{n}{k}=\binom{n}{k}+D_{t,n,k,s},$
where the second equality comes from Step 1 and the following equalities come from k+t2s=nk+t-2s=n.
Subcase 3.2.2: If t<n2t<n-2, then (t,n1,k1,s)(t,n-1,k-1,s) is valid and, hence, $r^{(1)}_{t,n,k,t-2s}=r(M^{(t,n-1)}_{k-1,k-1+t-2s})+\binom{n-1}{k}$ can be calculated via Eq 13 by induction. That is
@@
where the second equality comes from Eq 13 by induction, and the third equality comes from k+t2s=nk+t-2s=n.
From above, combining Eq 12, we can verify that Eq 13 holds when k+t2s=nk+t-2s=n.
Step 4: verify Eq 13 when $k>0$, k+t2s<nk+t-2s<n, 1<t<n11<t<n-1 and $s>0$.
@@
Note that (t1,n1,k1,s1)(t-1,n-1,k-1,s-1) is valid and, hence, we can calculate $r^{(2)}_{t,n,k,t-2s}$ via Eq 13 by induction.
Case 4.1.1: If ks<0k-s<0 or ks>ntk-s>n-t, then (k1)(s1)<0(k-1)-(s-1)<0 or (k1)(s1)>(n1)(t1)(k-1)-(s-1)>(n-1)-(t-1).
@@
$\qquad+\sum\limits_{m=t-3s+1+2k}^{n-s}\left(\binom{m-1}{k+t-2s-1}-\binom{m-1}{k-s-1}\right)+\sum\limits_{m=n-s+1}^{n}\binom{m-1}{k+t-2s-1}$
@@
First, it is easy to verify the inequality when t=n2t=n-2, then we assume t<n2t<n-2 in the following.
Case 4.2.1: If (k1)s<0(k-1)-s<0, then $(k-1)-(s-1)=k-s\leq 0$ and
@@
Case 4.2.2: If $0\leq(k-1)-s<\lfloor((n-1)-t)/2\rfloor$, then $0\leq 1\leq k-s\leq\lfloor(n-t-1)/2\rfloor\leq\lfloor(n-t)/2\rfloor$ and noting that $t-3s+2k-1\leq t-s-1+(n-t-1)\leq n-s-2$
$A^{(1)}_{t,n-1,k-1,s}-A^{(1)}_{t-1,n-1,k-1,s-1}=-\left(\binom{t+2k-3s-2}{t+k-2s-2}-\binom{t+2k-3s-2}{k-s-2}\right)$,
@@
$=$ $(B_{t,n-1,k-1,s}-B_{t-1,n-1,k-1,s-1})+(C_{t,n-1,k-1,s}-C_{t-1,n-1,k-1,s-1})-\binom{n-1}{k-1+t-2s}$
$=$ $C_{t,n-1,k-1,s}-C_{t-1,n-1,k-1,s-1}-\binom{n-1}{k-1+t-2s}$
$\leq$ $\binom{n-2}{k+t-2s-2}-\binom{n-1}{k-1+t-2s}-\sum_{m=n-s}^{n-2}\binom{m-1}{k+t-2s-1}$
$\leq$ $0.$
Case 4.2.4: If (k1)s>(n1)t(k-1)-s>(n-1)-t, then ks>ntk-s>n-t.